AB=BA=I and AC=CA=I. Then what about B an C
a. They are equal
b. they are not equal
c. Each is equal to A
d. None

Answer by Boreal(15235)

AB=I; B=A(inv)
BA=I; A=B(inv)
A=C (inv)
C=A (inv)
B=A (inv) and C=A(inv)
B and C are equal; they are equal to A(inv) but NOT to A.